Calculate each Ka using the information below.

Acid Used: KHP

Molar Mass of Acid: 204.22

Amount of KHP used: 1.04 g

Initial (mL) 0 2 4 6 8 10
Final (mL) 2 4 6 8 10 10
Amount Added (mL) 2 2 2 2 2 -
pH 4.24 4.27 4.30 4.46 4.51 4.63
Ka
Answer #1

Amount KHP Used = 1049 no o moles of KHP = 1.04 mole 204.22 vol. KHP Used for each case - 2 m2 = 0.0022 IKHPI = 1.ou. zou.R2for PH= 4: 30 pka= 2 4.30 + 0.4059 = iso 59 ka- antilog (gig059) = 9,855 xlo-lo for PH= 4.45 pkg= 24.46 + 0.4059 = 9:3259 ka